# Approximation in Hölder spaces

### v2

@@ Abstract.
For a modulus of continuity $\omega$ and normed spaces $X,Y$ we introduce and study the subspaces $\dot{\operatorname{VC}}^{0,\omega}_{\Gamma}(X,Y)$ of vanishing scales $\Gamma\in\{\operatorname{small},\operatorname{large},\operatorname{far}\}$ of the homogeneous Hölder space $\dot{C}^{0,\omega}(X,Y).$ For many couples $X$ and $Y$, we characterize the subspaces that admit approximations by smooth, Lipschitz and boundedly supported functions in terms of these three vanishing scales. In the particular case $X=\mathbb{R}^{n},$ we identify $\dot{\operatorname{VC}}^{0,\omega}_{\Gamma}(\mathbb{R}^{n},Y)=\mathrm{VMO}^{\omega}_{\Gamma}(\mathbb{R}^{n},Y)$ with the corresponding vanishing mean oscillation spaces, thus transferring these approximations also to these spaces.
Key words: smooth approximation, vanishing mean oscillation, Hölder spaces, Banach spaces
2020 Mathematics Subject Classification: 46T20, 41A29, 42B35, 46B20, 26B35
@@ 1. Introduction and main results
We consider normed spaces $X=(X,\|\cdot\|_{X}),$ $Y=(Y,\|\cdot\|_{Y})$ and moduli of continuity $\omega:(0,\infty)\to(0,\infty)$ and study the Hölder spaces $\dot{C}^{0,\omega}(X,Y)$ that consist of those functions $f$ for which
@@
These spaces are subspaces of uniformly continuous functions, in particular. We establish various full characterizations of approximability in the class $\dot{C}^{0,\omega}(X,Y)$ by Lipschitz and/or smooth and/or boundedly supported test functions. This is done in terms of the three vanishing scales that we define below, but before that let us make a few observations. It is easy to see that, if $Y$ a Banach space, then $\dot{C}^{0,\omega}(X,Y)$ also becomes a Banach space when equipped with the norm
@@
By dropping the term $\|f(0)\|_{Y},$ the seminorm $\|\cdot\|_{\dot{C}^{0,\omega}(X,Y)}$ becomes a norm when the functions $f\in\dot{C}^{0,\omega}(X,Y)$ are defined modulo constant equivalence classes. Moreover, in the particular case $\omega(t)=t^{\alpha},$ $\alpha\in(0,1),$ we recover the Hölder spaces $\dot{C}^{0,\alpha}(X,Y).$ Now, let us notate
@@
and then we define the three vanishing scales
$\dot{\operatorname{V}}^{0,\omega}_{\operatorname{small}}(X,Y)=\Big{\{}f:X\to Y\,:\,\lim_{\delta\to 0}\operatorname{osc}^{\omega}_{\delta}(f)=0\Big{\}},\,$
$\dot{\operatorname{V}}^{0,\omega}_{\operatorname{large}}(X,Y)=\Big{\{}f:X\to Y\,:\,\lim_{\delta\to\infty}\operatorname{osc}^{\omega}_{\delta}(f)=0\Big{\}},$
$\dot{\operatorname{V}}^{0,\omega}_{\operatorname{far}}(X,Y)=\Big{\{}f:X\to Y\,:\,\lim_{\delta\to\infty}\sup_{\min(\|x\|,\|y\|)>\delta}\operatorname{osc}^{\omega}_{(x,y)}(f)=0\Big{\}}.$
Then, we define the vanishing subspaces of $\dot{C}^{0,\omega}(X,Y).$
@@
For each scale $\Gamma\in\{\operatorname{small},\operatorname{large},\operatorname{far}\},$ let
@@
Whenever no confusion is possible, we drop the source and target spaces in the norms and simplify notation with $\|x\|_{X}=\|x\|$ and $\|f(x)\|_{Y}=|f(x)|;$ we also write $\dot{C}^{0,\omega}(X,Y)=\dot{C}^{0,\omega}(X)=\dot{C}^{0,\omega}.$
Remark 1.2.
A few remarks concerning the above definitions.
  • For $\Gamma\in\{\operatorname{small},\operatorname{large},\operatorname{far}\},$ the set $\dot{\operatorname{VC}}^{0,\omega}_{\Gamma}$ is closed with respect to the $\dot{C}^{0,\omega}$-seminorm. In particular, $(\dot{\operatorname{VC}}^{0,\omega}_{\Gamma},\|\cdot\|_{\dot{C}^{0,\omega}})$ are Banach spaces of modulo constant equivalence classes, provided the target space $Y$ is a Banach space.
  • By Lemma 2.1 at the beginning of Section 2 it follows that $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{far}}\subset\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{large}}.$ The converse inclusion is not true, as shown by elementary examples. It follows a posteriori that the natural scale $\Gamma=\operatorname{large}$ becomes artificial for some of our results, see e.g. Theorem 1.4 below. Moreover, in general $\dot{\operatorname{VC}}^{0,\omega}_{\Gamma}\not=\dot{\operatorname{V}}^{0,\omega}_{\Gamma}$ and $\dot{\operatorname{V}}^{0,\omega}_{\Gamma}\not=\dot{\operatorname{V}}^{0,\omega}_{\Gamma^{\prime}},$ for $\Gamma\not=\Gamma^{\prime}.$ Interestingly then, when we take the intersection of all the scales, we have $\dot{\operatorname{VC}}^{0,\omega}=\dot{\operatorname{V}}^{0,\omega}_{\operatorname{small}}\cap\dot{\operatorname{V}}^{0,\omega}_{\operatorname{far}}\cap\dot{\operatorname{V}}^{0,\omega}_{\operatorname{large}},$ and for the last inclusion ``""```$\supset$"` ` ⊃ ", we need the large scale on the right-hand side. These facts are not crucial to our results and we leave their verification to the interested reader.
Our modulus satisfies the following properties
@@
$\lim_{t\to 0}\frac{t}{\omega(t)}=0,$ (3)
and there exists a doubling constant $C_{\operatorname{db}}>1$ such that
$\omega(2t)\leq C_{\operatorname{db}}\omega(t),\qquad t\in(0,\infty).$ (4)
We mention explicitly whenever any of the modulus’ properties $(2),(3),(4)$ are used. Moreover, we always assume the modulus $\omega$ to be non-decreasing.
For the Hölder moduli $\alpha(t):=t^{\alpha},$ the classes $\dot{\operatorname{VC}}^{0,\alpha}_{\operatorname{small}}$ are sometimes called the little Hölder spaces. These are fundamental in the study of Lipschitz algebras in metric spaces, and we refer to the monograph by Weaver [Weaverbook, Chapters 4 and 8] for background. See also the recent monograph by D. Mitrea, I. Mitrea, and M. Mitrea [Mitreabook, Chapter 3] for a detailed exposition of the Hölder spaces in the setting of Ahlfors regular sets, their connection with functions of vanishing mean oscillation, and for an introduction to the little Hölder spaces.
@@ 1.1. Main results
During the last half a century, several fundamental results have been established in the theory of uniform approximation of continuous functions by smooth functions. The optimal order of smoothness arrangeable for these approximations is often determined by the regularity of a canonical or an equivalent renorming of the spaces $X$ and $Y.$ For background references on smoothness and renorming on Banach spaces, and the related approximation results, we refer to the monographs [BenyaminiLindenstrauss] by Benyamini and Lindenstrauss; [DevilleGodefroyZizler] by Deville, Godefroy, and Zizler; and [HajJoh10] by Hájek and Johanis. Closely related to our resuls are the articles by Azagra, Fry, and Keener [AzFrKe12], Cepedello-Boiso [Cepedello], Fry [Fry07], Hájek and Johanis [HajJoh09, HajJoh10], Lasry and Lions [LasryLions], Moulis [Moulis] and Pisier [Pisier]. We next clarify that uniform convergence and $\dot{C}^{0,\omega}$ convergence do not follow from each other.
Remark 1.3.
Endow $C^{0,\omega}(X,Y)$ with the norm $|||\cdot|||_{C^{0,\omega}}$ defined in (1). The next claims can be verified either directly from the definition or by elementary one-dimensional examples.
  • The convergence with respect to the norm $|||\cdot|||_{C^{0,\omega}}$ implies uniform convergence on bounded subsets of $X.$ But, in general, this convergence is not globally uniform on $X$, even when the sequence of functions involved is uniformly bounded on $X$.
  • The converse to claim (a) is not true: uniform convergence does not imply convergence with respect to the norm $|||\cdot|||_{C^{0,\omega}},$ even if one only looks at bounded subsets of $X$ for the norm $|||\cdot|||_{C^{0,\omega}},$ and even when the sequence of functions is uniformly bounded on $X.$
Our main result is the recognition of the scales $\{\operatorname{small},\operatorname{large},\operatorname{far}\}$ as precisely the concept that allows a complete description of the closure of smooth, Lipschitz and/or boundedly supported functions with respect to the $\dot{C}^{0,\omega}$ seminorm. In order for the reader to familiarize with the conclusions of our theorems, we begin by stating a result (without smoothness) on approximations in $\dot{\operatorname{VC}}^{0,\omega}(X,Y)$ by $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)$ functions with bounded support.
@@
Here and below, for any class of functions $\mathcal{C}(X,Y)\subset\dot{C}^{0,\omega}(X,Y),$ we denote by $\overline{\mathcal{C}(X,Y)}^{\dot{C}^{0,\omega}(X,Y)}$ the closure of $\mathcal{C}(X,Y)$ under the $\dot{C}^{0,\omega}(X,Y)$ seminorm. Naturally, here we refer to the family of those $f\in\dot{C}^{0,\omega}(X,Y)$ for which there exists a sequence $\{f_{n}\}_{n}\subset\mathcal{C}(X,Y)$ so that $\lim_{n}\|f_{n}-f\|_{\dot{C}^{0,\omega}(X,Y)}=0.$ The subscripted set $\mathcal{C}_{\operatorname{bs}}(X,Y)$ consists of those boundedly supported $h:X\to Y$ in the class $\mathcal{C}.$ In Theorem 1.4 the class $C_{\operatorname{bs}}(X,Y)$ denotes the space of continuous functions $f:X\to Y$ with bounded support, in particular. Naturally, the functions in $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)$ are continuous, and the intersection with $C_{\operatorname{bs}}(X,Y)$ indicates that the approximations have bounded support.
While Theorem 1.4 does not provide any smoothness in the approximations, it turns out to be a crucial step for the rest of our approximation theorems. For instance, letting $X=\mathbb{R}^{n}$ and $Y$ be a Banach space, we combine Theorem 1.4 with elementary mollification arguments to obtain $C^{\infty}_{\operatorname{bs}}(\mathbb{R}^{n},Y)$ approximations in Theorem 1.8 below and both theorems will be proved in Section 2.
Furthermore, our results are not confined to functions with a finite-dimensional source $X,$ and letting $X$ be a normed space and $Y=\mathbb{R},$ we prove smooth and Lipschitz approximation for the classes $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,\mathbb{R})$ and $\dot{\operatorname{VC}}^{0,\omega}(X,\mathbb{R}).$ (The corresponding approximations for $\mathbb{C}^{n}$-valued functions, in place of $\mathbb{R}$-valued, follow on each occassion after splitting them into their $n$-coordinates and then each to their real and imaginary parts.) In particular, with $X$ being a separable normed space, we have the following general result.
@@
Notice that, together with the $C^{k}$ order of smoothness, Theorem 1.5 guarantees Lipschitz approximations. Any separable normed space $X$ with an equivalent norm of class $C^{k}$ satisfies the assumption of Theorem 1.5, see Remark 3.6. In the particular case where $X$ is a separable Hilbert space, one can arrange $C^{\infty}$ smooth approximations, thus obtaining an infinite dimensional version of Theorem 1.8, see Corollary 3.7. We will restate Theorem 1.5 and prove it in Subsection 3.1.
Another class of spaces $X$ on which we obtain smooth approximations are super-reflexive Banach spaces. A theorem of Pisier [Pisier] says that any super-reflexive space admits a renorming with modulus of smoothness of power type $1+\alpha,$ for some $\alpha\in(0,1].$ See the property (39) below for the precise formulation. This class of spaces includes the Hilbert (separable or not) spaces, together with many of the classical Banach function spaces, such as the $L^{p}$ spaces, with $1<p<\infty;$ see Remark 3.11. The following theorem establishes $C^{1,\alpha}$ and Lipschitz approximations for functions of the classes $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X)$ and $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X).$
@@
Let $X$ be a super-reflexive space, and $\alpha\in(0,1]$ such that $X$ admits a renorming with modulus of smoothness of power type $1+\alpha.$ Then, for a modulus $\omega$ satisfying (2), (3) and (4), the following hold:
@@
This theorem will be restated and proved in Subsection 3.3, see Theorem 3.12. There, we will also show that the approximations can be taken to be Lipschitz in the case $Y=\mathbb{R}.$
1.2. Applications on $X=\mathbb{R}^{n}$
We provide a brief discussion of how our results relate to compactness of commutators of singular integral operators (SIOs). Compact operators are in particular bounded, and for many non-degenerate bounded SIOs $T$ and $p,q\in(1,\infty)$ there holds that
$\|[b,T]\|_{L^{p}(\mathbb{R}^{n})\to L^{q}(\mathbb{R}^{n})}\eqsim\|b\|_{X^{p,q}(\mathbb{R}^{n})},\quad X^{p,q}(\mathbb{R}^{n})=\begin{cases}\dot{C}^{0,\alpha(p,q)}(\mathbb{R}^{n}),&q>p,\\ \mathrm{BMO}(\mathbb{R}^{n})&q=p,\\ \dot{L}^{r(p,q)}(\mathbb{R}^{n}),&q<p,\end{cases}$ (5)
where $[b,T]f=bTf-T(bf)$ is the commutator and $b:\mathbb{R}^{n}\to\mathbb{C}$ is called the symbol of the commutator. The appearing exponents are defined through $\alpha(p,q)=n(1/p-1/q)$ and $1/q=1/r(p,q)+1/p,$ and $\dot{C}^{0,\alpha(p,q)}(\mathbb{R}^{n})$ is the inhomogeneous Hölder space, $\mathrm{BMO}(\mathbb{R}^{n})$ stands for the space of bounded mean oscillations and $\dot{L}^{r(p,q)}(\mathbb{R}^{n})$ is $L^{r(p,q)}(\mathbb{R}^{n})$ modulo additive constants. Emphasizing the recognition of the correct conditions on the symbol, the bound (5) is attributable to Coifman, Rochberg and Weiss [CRW], when $q=p;$ to Janson [Jan1978], when $q>p;$ and to Hytönen [HyLpLq], when $q<p,$ to where we also refer for a history leading to (5).

### v3

@@ Abstract.
We introduce new vanishing subspaces of the homogeneous Hölder space $\dot{C}^{0,\omega}(X,Y),$ in the generality of a doubling modulus $\omega$ and normed spaces $X$ and $Y.$ For many couples $X,Y,$ we show these vanishing subspaces to completely characterize those Hölder functions that admit approximations, in the Hölder seminorm, by smooth, Lipschitz and boundedly supported functions. Beyond the intrinsic interest of these results, we also present connections to bi-parameter harmonic analysis on the Euclidean space and in particular to compactness of the bi-commutator.
Key words and phrases:
smooth approximation, bounded support approximation, vanishing oscillation, Hölder space, Banach space
2020 Mathematics Subject Classification:
46T20, 41A29, 42B35, 46B20, 26B35
@@ 1. Introduction and main results
1.1. Introduction
We consider normed spaces $X,Y$ and moduli of continuity $\omega$ and study the Hölder spaces $\dot{C}^{0,\omega}(X,Y)$ that consist of those functions $f:X\to Y$ for which
@@ 1.1. Introduction
We establish various full characterizations of approximability in the class $\dot{C}^{0,\omega}(X,Y)$ by Lipschitz, smooth and boundedly supported functions in terms of three vanishing subspaces of $\dot{C}^{0,\omega}(X,Y)$ that demand vanishing behaviour of the Hölder norm of functions on small, large and far away scales. We will define these subspaces soon, but before that we make a few general observations.
First of all, if $Y$ is a Banach space, then so is $\dot{C}^{0,\omega}(X,Y),$ when equipped with the norm
@@
Dropping the term $\|f(0)\|_{Y},$ the seminorm $\|\cdot\|_{\dot{C}^{0,\omega}(X,Y)}$ becomes a norm when the functions are identified with equivalence classes modulo additive constants. In the particular case $\omega(t)=t^{\alpha},$ $\alpha\in(0,1),$ we recover the Hölder spaces $\dot{C}^{0,\alpha}(X,Y).$ Moreover, provided $\omega$ is any reasonable continuity moduli – at least $\omega(t)\to 0,$ as $t\to 0$ – the spaces $\dot{C}^{0,\omega}(X,Y)$ become subspaces of uniformly continuous functions.
The theory of uniform approximation of continuous functions by smooth functions has been extensively studied and developed in the last decades. For instance, when $X$ is a Hilbert space, Lasry and Lions [LasryLions] showed that Lipschitz real-valued functions can be uniformly approximated by $C^{1,1}$ Lipschitz functions. If $X$ is moreover separable, then a theorem of Moulis [Moulis] allows these approximations to be upgraded to $C^{\infty}$ approximations. Later, Cepedello-Boiso [Cepedello] obtained several results concerning $C^{1}$ or $C^{1,\alpha}$ approximation of uniformly continuous functions in super-reflexive Banach spaces $X$ – see also our Theorem 3.10 below and the very recent work of Johanis [Jo24], that contribute to this line of research. Then, Hájek and Johanis [HajJoh10] proved that when $X$ is separable with a $C^{k}$ smooth and Lipschitz bump, and either $X$ or $Y$ is super-reflexive, then uniformly continuous mappings admit approximations by $C^{k}$ maps that are still uniformly continuous. Lastly, we mention the real-analytic approximations of Azagra, Fry, and Keener [AzFrKe12] when $X$ is a Banach space with a separating polynomial; then, Lipschitz real-valued functions on $X$ can be uniformly approximated by real analytic Lipschitz functions. For more background references on smoothness and renorming on Banach spaces, and the related approximation results, we refer to the monographs [BenyaminiLindenstrauss], by Benyamini and Lindenstrauss; [DevilleGodefroyZizler], by Deville, Godefroy, and Zizler; and [HajJoh10], by Hájek and Johanis.
In the present article, we prove full characterizations of approximation with Hölder type rate of convergence taking the place of uniform convergence. We note that these two types of convergence do not follow from each other; see Appendix A for a detailed comparison with simple observations and examples. Our main result is best described as the recognition of the vanishing scales small, large and far (below) as precisely the concept that allows a complete description of the closure of smooth, Lipschitz and boundedly supported functions with respect to the $\dot{C}^{0,\omega}$ seminorm. We notate
@@
and then define these three vanishing scales as
$\dot{\operatorname{V}}^{0,\omega}_{\operatorname{small}}(X,Y)$ =$=\big{\{}f:X\to Y\,:\,\lim_{\delta\to 0}\operatorname{osc}^{\omega}_{\delta}(f)=0\big{\}},\,$
$\dot{\operatorname{V}}^{0,\omega}_{\operatorname{large}}(X,Y)$ =$=\big{\{}f:X\to Y\,:\,\lim_{\delta\to\infty}\operatorname{osc}^{\omega}_{\delta}(f)=0\big{\}},$
$\dot{\operatorname{V}}^{0,\omega}_{\operatorname{far}}(X,Y)$ =$=\big{\{}f:X\to Y\,:\,\lim_{\delta\to\infty}\sup_{\min(\|x\|,\|y\|)>\delta}\operatorname{osc}^{\omega}_{(x,y)}(f)=0\big{\}}.$
@@
For each scale $\Gamma\in\{\operatorname{small},\operatorname{large},\operatorname{far}\},$ set
@@
If $Y$ is a Banach space, then so are all of $\dot{\operatorname{VC}}^{0,\omega}_{\Gamma}(X,Y)$ as spaces of modulo additive constant equivalence classes; for full details, see Remark 2.1 below. From now on, whenever no confusion is possible, we start dropping the source and target spaces from the norms and simplify notation with $\|x\|_{X}=\|x\|$ and $\|f(x)\|_{Y}=|f(x)|;$ in the same way, we write $\dot{C}^{0,\omega}(X,Y)=\dot{C}^{0,\omega}(X)=\dot{C}^{0,\omega},$ and so on.
We assume our moduli $\omega$ to be non-decreasing and satisfy
@@
$\lim_{t\to 0}t/\omega(t)=0,$ (3)
$\omega(2t)\leq C_{\operatorname{db}}\omega(t),\qquad\mbox{for}\qquad t\in(0,\infty),$ (4)
where the doubling constant $C_{\operatorname{db}}>1.$ This is a big class of moduli that in particular covers the Hölder modulus $\alpha(t):=t^{\alpha},$ when $\alpha\in(0,1),$ but of course other more exotic moduli are also covered. The classes $\dot{\operatorname{VC}}^{0,\alpha}_{\operatorname{small}}$ are also called the little Hölder spaces. These are important in the study of Lipschitz algebras in metric spaces, and we refer to the monograph by Weaver [Weaverbook, Chapters 4 and 8] for a background. See also the recent monograph by D. Mitrea, I. Mitrea, and M. Mitrea [Mitreabook, Chapter 3] for a detailed exposition of the Hölder spaces in the setting of Ahlfors regular sets and their connection with functions of vanishing mean oscillation.
Notation
We denote $A\lesssim B$, if $A\leq CB$ for some constant $C>0$ depending only on an underlying function space that is fixed relative to ongoing considerations. For example, $C$ could depend on the modulus $\omega,$ or the dimension $n$ in the case $X=\mathbb{R}^{n}.$ Then $A\sim B$, if $A\lesssim B$ and $B\lesssim A.$ Subscripts or variables on constants and quantifiers $C_{a},C(a),\lesssim_{a}$ indicate their dependence on those subscripts.
@@ 1.2. Main results
We begin by stating a result without any smoothness on approximations by functions with bounded support. Many of our later results build on this one.
@@
Here and below, for any $\mathcal{C}(X,Y)\subset\dot{C}^{0,\omega}(X,Y),$ we denote by $\overline{\mathcal{C}(X,Y)}^{\dot{C}^{0,\omega}(X,Y)}$ the closure of $\mathcal{C}(X,Y)$ under the $\dot{C}^{0,\omega}(X,Y)$ seminorm. We refer here to the family of those $f\in\dot{C}^{0,\omega}(X,Y)$ for which there exists a sequence $\{f_{n}\}_{n}\subset\mathcal{C}(X,Y)$ so that $\lim_{n}\|f_{n}-f\|_{\dot{C}^{0,\omega}(X,Y)}=0.$ The subscripted set $\mathcal{C}_{\operatorname{bs}}(X,Y)$ consists of boundedly supported functions in $\mathcal{C}(X,Y).$ In Theorem 1.2 $C_{\operatorname{bs}}(X,Y)$ denotes continuous functions with bounded support. In particular, the functions in $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)$ are continuous, and the intersection with $C_{\operatorname{bs}}(X,Y)$ gives the approximations bounded support.
While Theorem 1.2 does not provide any smoothness, it turns out to be a crucial step in our other theorems. For instance, letting $X=\mathbb{R}^{n}$ and $Y$ be a Banach space, Theorem 1.2 coupled with a mollification argument gives $C^{\infty}_{\operatorname{bs}}(\mathbb{R}^{n},Y)$ approximations, this is Theorem 1.6 below. Both of these theorems will be proved in Section 2.
Our results are not confined to functions with a finite-dimensional source $X,$ and letting $Y=\mathbb{R},$ we prove smooth and Lipschitz approximation for the classes $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,\mathbb{R})$ and $\dot{\operatorname{VC}}^{0,\omega}(X,\mathbb{R}).$ (The corresponding approximations for $\mathbb{C}^{n}$-valued functions in place of $\mathbb{R}$-valued then follow after splitting into the coordinate functions and then each of these to the real and imaginary parts.) In particular, with $X$ being a separable normed space, we have the following general result.
@@
Notice that, together with the $C^{k}$ order of smoothness, Theorem 1.3 guarantees Lipschitz approximations. In particular, any separable normed space $X$ with an equivalent norm of class $C^{k}$ satisfies the assumption of Theorem 1.3, see Remark 3.6. Moreover, when $X$ is a separable Hilbert space, one can arrange $C^{\infty}$ smooth approximations, thus obtaining an infinite dimensional version of Theorem 1.6, see Corollary 3.7. We will restate Theorem 1.3 and prove it in Subsection 3.1.
Another class of spaces $X$ on which we obtain smooth approximations are super-reflexive Banach spaces. A theorem of Pisier [Pisier] says that any super-reflexive space admits a renorming with modulus of smoothness of power type $1+\alpha,$ for some $\alpha\in(0,1]$ – see (35) below for the precise formulation. These spaces contain the Hilbert (separable or not) spaces, together with many of the classical Banach function spaces, such as the $L^{p}$ spaces, with $1<p<\infty;$ see Remark 3.11. In the following theorem we establish $C^{1,\alpha}$ and Lipschitz approximations for the classes $\dot{\operatorname{VC}}^{0,\omega}(X)$ and $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X).$
@@
Let $X$ be a super-reflexive space that admits a renorming with modulus of smoothness of power type $1+\alpha,$ $\alpha\in(0,1].$ Then, for a modulus $\omega$ satisfying (2), (3) and (4), the following hold:
@@
Theorem 1.5 will be restated and proved in Subsection 3.3, see Theorem 3.12. There, we will also show that the approximations can be taken to be Lipschitz in the case $Y=\mathbb{R}.$
1.3. Applications on the Euclidean space
In the previous section we presented results that guarantee nice approximations, provided we know pointwise vanishing type conditions. When $X=\mathbb{R}^{n},$ we present formulations of these results in terms of mean oscillations type conditions, that are very convenient when studying the boundedness and compactness of commutators of Calderón-Zygmund operators (CZOs). In the one parameter setting our results recover important components to the theory, but from a more general approach. In the bi-parameter setting our results have very recently found applications by Martikainen and the second named author [MarOik24] in the study of the compactness of the bi-commutator. We describe all of this in full detail below, but to not get too far ahead of ourselves, let us begin by recapping boundedness and compactness characterizations of commutators in the one-parameter setting.
For a non-degenerate CZO $T$ on $\mathbb{R}^{n},$ a function $b\in L^{1}_{\operatorname*{loc}}(\mathbb{R}^{n},\mathbb{C})$ (symbol of the commutator) and exponents $p,q\in(1,\infty),$ the commutator has the following mapping properties
$\|[b,T]\|_{L^{p}(\mathbb{R}^{n})\to L^{q}(\mathbb{R}^{n})}\sim\|b\|_{X^{p,q}(\mathbb{R}^{n})},$ (5)
where the space $X^{p,q}(\mathbb{R}^{n})$ is determined as follows. Given $p,q\in(1,\infty),$ denote by $\alpha(p,q)$ and $r(p,q)$ the exponents uniquely determined through the relations $\alpha(p,q)/n=1/p-1/q$ and $1/q=1/r(p,q)+1/p.$ Then, $X^{p,q}(\mathbb{R}^{n})$ is the space of bounded mean oscillations $\mathrm{BMO}(\mathbb{R}^{n}),$ when $p=q;$ the fractional BMO space $\mathrm{BMO}^{\alpha(p,q)}(\mathbb{R}^{n}),$ when $p<q;$ and the Lebesgue space $L^{r(p,q)}(\mathbb{R}^{n})$ modulo additive constants $\dot{L}^{r(p,q)}(\mathbb{R}^{n}),$ when $p>q.$ (The reader can recover the definitions of the spaces $\mathrm{BMO},\mathrm{BMO}^{\alpha}$ from Definition 1.7 below with the ``$\mbox{moduli}"$ $t\mapsto t^{0}$ and $t\mapsto t^{\alpha},$ respectively.) Emphasizing the recognition of the correct function spaces $X^{p,q}(\mathbb{R}^{n})$ to which the symbol belongs, the result (5) is due to Nehari [Nehari1957] ($n=1$) and Coifman, Rochberg, Weiss [CRW] ($n\geq 2$), when $q=p;$ due to Janson [Jan1978], when $q>p;$ and due to Hytönen [HyLpLq], when $q<p.$ We also refer the reader to the introductory section of [HyLpLq] for precise definitions and a fuller history of what was outlined above. Next, we describe the corresponding off-diagonal compactness characterization.